OpenStudy (anonymous):

x^2-2x-24=0 SHOW WORK SOLVE FOR VARIABLE

OpenStudy (wasiqss):

(x-6)(x+4)

OpenStudy (mysesshou):

(x-6)(x+4) = 0 (x-6) = 0 and (x+4) = 0 x=6 and x= -4